OpenStudy (anonymous):

A skier with a mass of 60.0kg pushes off with a horizontal speed of 3.0 m/s from the top of a snow covered hill. As he slides down the hill, his height decreases by 42m over a distance of 100 m of travel. How much will his kinetic energy increase by the time he reaches the bottom of the hill (assuming no friction)?

OpenStudy (anonymous):

Total Energy Initially = KEi + PEi = 1/2 mv^2 + mgh = 1/2(60)(3)^2 + (60)(9.8)(42) = 24,966 J Total Energy must be conserved. Since the Total Energy Initially was 24,966 J and we assume that all energy of both kinetic and potential is converted to kinetic by the bottom of the hill, the Total Energy Initially is equal to the Kinetic Energy at the bottom of the hill. KE at bottom = 24,966 J

OpenStudy (anonymous):

Thank you so much :)